DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 23 Padelford C-36 3:45 PM

On conformally compact Einstein manifolds

Xiaodong Wang (Stanford University)

Conformally compact Einstein manifolds were first studied by mathematicians more than a decade ago. Recently the subject has been greatly revived by exciting developments in string theory. New insights as well as new problems have emerged. In my talk I will present some results on the subject, focusing on the interaction between the global geometry of the interior and the conformal geometry on the boundary.